# Nes2Net: A Lightweight Nested Architecture for Foundation Model Driven Speech Anti-spoofing

### v1

@@ Abstract
Speech foundation models have significantly advanced various speech-related tasks by providing exceptional representation capabilities. However, their high-dimensional output features often create a mismatch with downstream task models, which typically require lower-dimensional inputs. A common solution is to apply a dimensionality reduction (DR) layer, but this approach increases parameter overhead, computational costs, and risks losing valuable information. To address these issues, we propose Nested Res2Net (Nes2Net), a lightweight back-end architecture designed to directly process high-dimensional features without DR layers. The nested structure enhances multi-scale feature extraction, improves feature interaction, and preserves high-dimensional information. We first validate Nes2Net on CtrSVDD, a singing voice deepfake detection dataset, and report a 22% performance improvement and an 87% back-end computational cost reduction over the state-of-the-art baseline. Additionally, extensive testing across four diverse datasets—ASVspoof 2021, ASVspoof 5, PartialSpoof, and In-the-Wild—covering fully spoofed speech, adversarial attacks, partial spoofing, and real-world scenarios, consistently highlights Nes2Net’s superior robustness and generalization capabilities. The code package and pre-trained models are available at https://github.com/Liu-Tianchi/Nes2Net.
Index Terms:
DeepFake detection, speech anti-spoofing, Res2Net, Nes2Net, SSL, speech foundation model
@@ I Introduction
While the speech foundation models offer exceptional representations, their high-dimensional feature outputs present significant challenges for downstream tasks. Downstream models used in tasks like speech anti-spoofing typically require lower-dimensional features [15, 16, 17]. To address this mismatch, a common approach is to introduce a dimensionality reduction (DR) layer, usually implemented as a fully connected (FC) layer for transforming high-dimensional features into lower-dimensional features. However, this conventional strategy presents notable drawbacks. Since downstream classifiers are usually compact, the DR layer alone often consumes a substantial portion of the parameters and computational resources within the entire back-end model. Moreover, directly projecting high-dimensional features in a one-shot manner through an FC layer leads to the loss of important information, reducing the effectiveness of speech foundation models. These issues highlight the need for a more efficient and effective solution to bridge the dimensionality gap and fully utilize speech foundation models in downstream tasks.
@@
  • Enhanced Performance, Efficiency, and Generalization: Our method demonstrates significant improvements in efficiency, achieving performance improvements of 22% while reducing computational costs by 87% compared to the SOTA baseline on the CtrSVDD dataset. Further experiments conducted on four additional datasets across various scenarios demonstrate strong generalization capability and consistently superior performance.
@@ II-A Res2Net
Res2Net [18] is a well-known architecture designed to extract multi-scale features. Unlike ResNet [19], Res2Net uses hierarchical residual connections within a single block, allowing it to capture patterns across varying receptive fields simultaneously [18]. This design offers proven advantages in speech-related tasks, such as SV [20, 21, 22] and anti-spoofing [23, 24, 25], where capturing subtle variations and complex acoustic patterns is important. As shown in Fig. 1, Res2Net (highlighted using a light red block) can also serve as a classifier within a speech foundation model-based anti-spoofing system. Its ability to extract multi-scale features has led to superior performance to conventional models, that motivates the design of Nested Res2Net in this work.
II-B Traditional Speech Anti-Spoofing Models
Traditional speech anti-spoofing models have evolved to effectively detect speech deepfakes [26, 27]. For instance, the Channel-wise Gated Res2Net (CG-Res2Net) [23] introduces a gating mechanism within the Res2Net architecture, enabling dynamic selection of channel-wise features to enhance generalization to unseen attacks. A widely recognized model among traditional approaches is AASIST [26], which employs spectro-temporal graph attention layers to capture both temporal and spectral artifacts, thereby achieving efficient and accurate detection. Given AASIST’s SOTA performance and its wide adoption in recent anti-spoofing challenges [16, 28], we consider it as our main baseline for evaluation.
Although traditional anti-spoofing models are characterized by their compact size and rapid inference capabilities, they typically rely on hand-crafted acoustic features, which limit their adaptability to diverse spoofing methods. With the rise of deepfakes and adversarial attacks, stronger and more flexible representations are needed.
@@ II-C Speech Foundation Models
Speech foundation models, often referred to as Self-Supervised Learning (SSL) models due to their typical pre-training on large amounts of unlabeled speech data using self-supervised learning techniques. Examples include wav2vec 2.0 [1], HuBERT [2], and WavLM [3]. Unlike traditional acoustic feature-based approaches, which rely on hand-crafted features and are often task-specific, SSL models learn rich and generalized speech representations that can be effectively adapted to various downstream applications. This allows them to achieve superior performance in speech-related tasks, including speech anti-spoofing.
@@ II-D Speech Foundation Model-based Anti-spoofing
As discussed in the previous subsection, speech foundation models can capture more informative representations than traditional handcrafted or raw acoustic features [3]. This makes them highly effective for speech anti-spoofing, as they generalize well across datasets and are more robust to unseen attacks [15]. As a result, many recent anti-spoofing systems increasingly adopt these models as front-ends, feeding their features to back-end classifiers and consistently outperforming traditional models [29, 16, 30].
@@
Following the aggregation layer, the resulting features are passed to the back-end classifier, as shown in the green box of Fig. 1. Existing methods typically use a DR layer, which reduces the high-dimensional feature of channel $N$ (commonly $N=1024$ [1, 3, 32]) to a lower dimension $D$ (e.g., $D=128$ [15, 16] or $D=144$ [17, 33]) to match the classifier model’s input requirements. The classifier model then extracts features from the DR layer outputs and produces the final score. As illustrated in the red box of Fig. 1, commonly used classifier structures include traditional models such as ResNet [19], Res2Net [18], ECAPA-TDNN [34], and AASIST [26].
@@ II-E Limitation of Dimensionality Reduction Layer
Across various back-end models, the DR layer, despite being just a single layer, consistently accounts for a substantial share of parameters and computational cost, underscoring its resource-intensive nature. For instance, the DR layer accounts for 21% to 29% of the parameters across ResNet, Res2Net, ECAPA, and AASIST. In terms of computational cost, the DR layer generally contributes at least one-third of the total MACs. AASIST is the only exception, where the DR layer accounts for just 4% of the MACs, primarily because the overall MACs of AASIST are significantly larger—an order of magnitude higher than those of other models.
@@ III-A Proposed Nested Res2Net (Nes2Net)
Efficiency and Retention of Rich Feature Information: The analysis in Section II-E reveals the limitations of employing the DR layer, highlighting its significant computational overhead and potential information loss. These drawbacks motivate the need for a more efficient solution—one that bridges the dimensionality gap without sacrificing the richness of high-dimensional representations from speech foundation models. To this end, Nes2Net entirely removes the DR layer, directly processing high-dimensional features to retain their intrinsic richness and minimize unnecessary computational costs. By bypassing the DR layer, Nes2Net prevents the information bottleneck typically caused by early dimensionality reduction. This ensures the preservation of detailed representations essential for accurately distinguishing genuine speech from spoofed audio. Furthermore, the nested structure optimizes computational efficiency by systematically refining and managing feature representations across multiple groups, ensuring maximum utilization of high-dimensional features with minimal computational overhead.
@@
Enhanced Multi-Scale Feature Interaction and Expressiveness: While the Res2Net architecture effectively extracts multi-scale features through hierarchical splits, it exhibits significant limitations when processing high-dimensional features directly, especially with large split scales $s$. Specifically, Res2Net suffers from feature dilution [18], redundant transformations [43], and restricted interactions among channels. Excessive splitting fragments the features, weakening their expressiveness, and repetitive transformations increase computational redundancy, potentially causing overfitting. Moreover, closely related information can be distributed across distant subsets, limiting effective cross-channel interactions.
@@
To overcome these limitations, as illustrated in Fig. 1, we propose a novel Nested Res2Net (Nes2Net) architecture that introduces a hierarchical nesting structure. This added degree of flexibility significantly enhances the model’s representational power. Each nested layer progressively refines features by building upon outputs from preceding layers and also incorporates efficient local cross-channel attention mechanisms [44, 45], strengthening interactions across channels. This holistic feature extraction approach enables Nes2Net to comprehensively capture intricate speech patterns. Moreover, the cumulative refinement effectively mitigates the issue of feature dilution, preserving rich and expressive multi-scale information. Benefiting from the structural advantages of the nesting strategy, the need for excessive fine-grained splits is reduced, effectively mitigating redundant transformations. This approach also minimizes unnecessary computations, resulting in a compact yet highly expressive model.
@@
Critically, overfitting is a well-known challenge in speech anti-spoofing tasks, often leading to degraded performance in cross-domain scenarios. Previous studies [26, 23], particularly with compact models like AASIST and Res2Net (both with fewer than 500k parameters), have shown that smaller models can help reduce overfitting. Our experiments with these models confirm that simply increasing their size does not always lead to better performance and can, in fact, make overfitting worse. As a result, improving feature quality through smarter model structure design becomes more important than just scaling up the model. The nested architecture of Nes2Net provides clear benefits—it maintains computational efficiency while reducing the risk of overfitting. In addition, its improved structure produces more generalizable features, leading to stronger performance, especially in unseen and cross-domain scenarios.
@@ III-A1 Outer Layer
The outer layer of Nes2Net adopts a structure similar to that of Res2Net. The high-dimensional features produced by the speech foundation model are uniformly split into $s_{1}$ feature map subsets, denoted by $x_{i}$, where $i\in\{1,2,\ldots,s_{1}\}$. Each feature subset $x_{i}$ has the same spatial size but contains only $\frac{1}{s_{1}}$ of the channels of the input feature map. With the exception of $x_{1}$, each $x_{i}$ is paired with a corresponding nested layer, denoted by $\mathbf{K}_{i}()$. The output of $\mathbf{K}_{i}()$, represented as $y_{i}$, is computed as follows:
@@
where $x_{i}$ is first added to the output of $\mathbf{K}_{i-1}()$, and the resulting feature map is then fed into $\mathbf{K}_{i}()$ for further processing. All $y_{i}$ features are concatenated along the channel dimension. Due to the combinatorial explosion effect [18], the output features encapsulate a fusion of receptive field characteristics across different scales and frame levels. These features are then pooled along the time axis to convert frame-level features into utterance-level representations, which are subsequently used to compute the final classification score.
@@ III-A2 Nested Layer
The nested layer acts as the core module responsible for processing the outer layer’s intermediate features, denoted bys $x^{\prime}_{i}$, where $i\in\{2,\ldots,s_{1}\}$. Based on Eq. 1, $x^{\prime}_{i}$ is defined as:
@@
Each nested layer module $\mathbf{K}_{i}()$ is designed to extract multi-scale representations from its input while maintaining computational efficiency. As shonw in Fig. 1, the structure of $\mathbf{K}_{i}()$ follows a SE-Res2Net-like design, but its input is the subset feature $x^{\prime}_{i}$ from the outer layer of Nes2Net. Specifically, each nested layer consists of the following components:
@@
Multi-Scale Feature Extraction: To enable multi-scale processing, the input feature map $x^{\prime}_{i}$ is equally split into $s_{2}$ subsets along the channel dimension, denoted by $x_{i,j}$, where $j\in\{1,2,\ldots,s_{2}\}$. Each subset undergoes separate transformations through convolutional operations $\mathbf{M}_{j}$ with varying receptive fields, yielding $y_{i,j}$, formulated as:
${y}_{i,j}=\begin{cases}{x}_{i,j}&j=1;\\ \mathbf{M}_{j}({x}_{i,j}+{y}_{i,j-1})&1<j\leq s_{2}.\end{cases}$ (3)
@@
Residual Connections: To enhance gradient flow and stabilize training, a residual connection is applied by adding the input of $x^{\prime}_{i}$ to itss output $y_{i}$. This design preserves the original information while incorporating newly learned features.
In summary, the nested layer is lightweight and highly efficient. Its ability to capture multi-scale features ensures that the model maintains robustness and generalization across diverse conditions.
@@ III-B Enhanced Nested Res2Net (Nes2Net-X)
While Nes2Net efficiently addresses the high-dimensional feature issue, it relies on an additive combination method within the nested layer, which potentially limits the flexibility and effectiveness of the feature extraction by assigning equal importance to all features. To further enhance the representational capability of Nes2Net, we propose an improved variant, Nes2Net-X, that incorporates concatenation followed by a weighted summation within its nested layers. This design explicitly preserves feature subset individuality before fusion and employs learnable weights to adaptively combine these subsets, offering several advantages:
@@
The Nes2Net-X consists of the following components:
Feature Splitting and Processing: This component is the same as that in the Nes2Net nested layer. The input feature $x^{\prime}_{i}$ is equally split into $s_{2}$ subsets along the channel dimension, denoted by $x_{i,j}$, where $j\in\{1,2,\ldots,s_{2}\}$. Each subset $x_{i,j}$ undergoes a convolutional operation to extract feature representations.
Feature Concatenation: The outputs of the convolutional layer are denoted as $z_{i,j}$. In Nes2Net-X, instead of summing the processed features as in the original Nes2Net nested layer, each current subset $x_{i,j}$ is concatenated with the previous output $z_{i,j-1}$ before being processed.
$y_{i,j}=\sum_{k=1}^{s}w_{i,j,k}\cdot z_{i,j,k}$ (4)
where $s$ denotes the number of subsets, $w_{i,j,k}$ represents the weight for the $k$-th subset features $z_{i,j,k}$. The weighted summation allows the model to prioritize more informative features dynamically while suppressing less relevant ones, improving the overall robustness and performance of the nested layer.

### v2

# This information aligns with the guidelines available at:

@@ Abstract
Speech foundation models have significantly advanced various speech-related tasks by providing exceptional representation capabilities. However, their high-dimensional output features often create a mismatch with downstream task models, which typically require lower-dimensional inputs. A common solution is to apply a dimensionality reduction (DR) layer, but this approach increases parameter overhead, computational costs, and risks losing valuable information. To address these issues, we propose Nested Res2Net (Nes2Net), a lightweight back-end architecture designed to directly process high-dimensional features without DR layers. The nested structure enhances multi-scale feature extraction, improves feature interaction, and preserves high-dimensional information. We first validate Nes2Net on CtrSVDD, a singing voice deepfake detection dataset, and report a 22% performance improvement and an 87% back-end computational cost reduction over the state-of-the-art baseline. Additionally, extensive testing across four diverse datasets: ASVspoof 2021, ASVspoof 5, PartialSpoof, and In-the-Wild, covering fully spoofed speech, adversarial attacks, partial spoofing, and real-world scenarios, consistently highlights Nes2Net’s superior robustness and generalization capabilities. The code package and pre-trained models are available at https://github.com/Liu-Tianchi/Nes2Net.
@@ I Introduction
While speech foundation models offer exceptional representations, their high-dimensional feature outputs present significant challenges for downstream tasks. Downstream models used in tasks like speech anti-spoofing typically require lower-dimensional features [15, 16, 17]. To address this mismatch, a common approach is to introduce a dimensionality reduction (DR) layer, usually implemented as a fully connected (FC) layer for transforming high-dimensional features into lower-dimensional features. However, this conventional strategy presents notable drawbacks. Given that downstream classifiers are typically compact [15, 16], the DR layer alone often consumes a substantial portion of the parameters and computational resources within the entire back-end model. Moreover, directly projecting high-dimensional features in a one-shot manner through an FC layer leads to the loss of important information, reducing the effectiveness of speech foundation models. These issues highlight the need for a more efficient and effective solution to bridge the dimensionality gap and fully utilize speech foundation models in downstream tasks.
@@
  • Enhanced Performance, Efficiency, and Generalization: Our method demonstrates a 22% performance gain and an 87% reduction in computational costs compared to the state-of-the-art baselines on the CtrSVDD dataset. Further experiments conducted on four additional datasets across various scenarios demonstrate strong generalization capability and consistently superior performance.
@@ II-A Res2Net
Res2Net [18] is a well-known architecture designed to extract multi-scale features. Unlike ResNet [19], Res2Net uses hierarchical residual connections within a single block, allowing it to capture patterns across varying receptive fields simultaneously [18]. This design offers proven advantages in speech-related tasks, such as SV [20, 21, 22] and anti-spoofing [23, 24, 25], where capturing subtle variations and complex acoustic patterns is important. As shown in Fig. 1, Res2Net (highlighted using a light red block) can also serve as a classifier within a speech foundation model-based anti-spoofing system. Its ability to extract multi-scale features has led to superior performance over conventional models and motivates the design of Nested Res2Net in this work.
II-B Hand-crafted Feature-based Speech Anti-Spoofing Models
Hand-crafted acoustic features (such as MFCC) are common choices for many earlier speech anti-spoofing systems. These systems have evolved to effectively detect speech deepfakes [26, 27]. For instance, the Channel-wise Gated Res2Net (CG-Res2Net) [23] introduces a gating mechanism within the Res2Net architecture, enabling dynamic selection of channel-wise features to enhance generalization to unseen attacks. A widely recognized model is AASIST [26], which employs spectro-temporal graph attention layers to capture both temporal and spectral artifacts, thereby achieving efficient and accurate detection. Given AASIST’s SOTA performance and its wide adoption in recent anti-spoofing challenges [16, 28], we consider it as our main baseline for evaluation.
@@ II-C Speech Foundation Models
Speech foundation models are often referred to as Self-Supervised Learning (SSL) models due to their typical pre-training on large amounts of unlabeled speech data using self-supervised learning techniques. Examples include wav2vec 2.0 [1], HuBERT [2], and WavLM [3]. Unlike hand-crafted acoustic features, which are limited in their ability to adapt to diverse and complex conditions, self-supervised learning (SSL) models learn rich and generalized speech representations that can be effectively adapted to various downstream applications. This allows them to achieve superior performance in speech-related tasks, including speech anti-spoofing.
@@ II-D Speech Foundation Model-based Anti-spoofing
As discussed in the previous subsection, speech foundation models can capture more informative representations than handcrafted or raw acoustic features [3]. This makes them highly effective for speech anti-spoofing, as they generalize well across datasets and are more robust to unseen attacks [15]. As a result, many recent anti-spoofing systems increasingly adopt these models as front-ends, feeding their features to the back-end classifiers and consistently outperforming traditional models [29, 16, 30].
@@
Following the aggregation layer, the resulting features are passed to the back-end classifier, as shown in the green box of Fig. 1. Existing methods typically use a DR layer, which reduces the high-dimensional features of $N$ channels (commonly $N=1024$ [1, 3, 32]) to a lower dimension $D$ (e.g., $D=128$ [15, 16] or $D=144$ [17, 33]) to match the classifier’s input requirements. The classifier model then extracts features from the DR layer outputs and produces the final score. As illustrated in the red box of Fig. 1, commonly used classifier structures include traditional models such as ResNet [19], Res2Net [18], ECAPA-TDNN [34], and AASIST [26].
@@ II-E Limitation of Dimensionality Reduction Layer
Across various back-end models, the DR layer, despite being just a single layer, consistently accounts for a substantial share of parameters and computational cost, underscoring its resource-intensive nature. For instance, the DR layer accounts for 21% to 29% of the parameters across ResNet, Res2Net, ECAPA, and AASIST. In terms of computational cost, the DR layer generally contributes at least one-third of the total MACs. AASIST is the only exception, where the DR layer accounts for just 4% of the MACs, primarily because its overall MAC count is an order of magnitude higher than that of other models.
@@ III-A Proposed Nested Res2Net (Nes2Net)
Efficiency and Retention of Rich Feature Information: The analysis in Section II-E reveals the limitations of employing the DR layer. Building upon the observations, Nes2Net entirely removes the DR layer, directly processing high-dimensional features to retain their intrinsic richness and minimize unnecessary computational costs. By bypassing the DR layer, Nes2Net prevents the information bottleneck typically caused by early dimensionality reduction. This ensures the preservation of detailed representations essential for accurately distinguishing genuine speech from spoofed audio.
@@
Enhanced Multi-Scale Feature Interaction and Expressiveness: While the Res2Net architecture effectively extracts multi-scale features through hierarchical splits, it exhibits significant limitations when processing high-dimensional features directly, especially with large split scales $s$. Specifically, Res2Net suffers from feature dilution [18], redundant transformations [43], and restricted interactions among channels. Excessive splitting fragments the features, weakening their expressiveness, and repetitive transformations increase computational redundancy, potentially causing overfitting. Moreover, closely related information can be distributed across non-adjacent subsets, limiting effective cross-channel interactions.
@@
To overcome these limitations, as illustrated in Fig. 1, we propose a novel Nested Res2Net (Nes2Net) architecture that introduces a hierarchical nesting structure. This additional degree of flexibility significantly enhances the model’s representational capability. Each nested layer progressively refines features by building upon outputs from preceding layers and also incorporates efficient local cross-channel attention mechanisms [44, 45], strengthening interactions across channels. This holistic feature extraction approach enables Nes2Net to comprehensively capture intricate speech patterns. Moreover, the cumulative refinement effectively mitigates the issue of feature dilution, preserving rich and expressive multi-scale information. Benefiting from the structural advantages of the nesting strategy, the need for excessive fine-grained splits is reduced, effectively mitigating redundant transformations. This approach also minimizes unnecessary computations, resulting in a compact yet highly expressive model.
@@
Critically, overfitting is a well-known challenge in speech anti-spoofing tasks, often leading to degraded performance in cross-domain scenarios. Previous studies [26, 23], particularly with compact models like AASIST and Res2Net (both with fewer than 500k parameters), have shown that smaller models can help reduce overfitting. Our experiments with these models confirm that simply increasing their size does not always lead to better performance and can, in fact, make overfitting worse. As a result, improving feature quality through smarter model structure design becomes more important than just scaling up the model. The nested architecture of Nes2Net provides clear benefits as it maintains computational efficiency while reducing the risk of overfitting.
@@ III-A1 Outer Layer
The outer layer of Nes2Net adopts a structure similar to that of Res2Net. The high-dimensional features produced by a speech foundation model are uniformly split into $s_{1}$ feature map subsets, denoted by $x_{i}$, where $i\in\{1,2,\ldots,s_{1}\}$. Each feature subset $x_{i}$ has the same spatial size but contains only $\frac{1}{s_{1}}$ of the channels of the input feature map. With the exception of $x_{1}$, each $x_{i}$ is paired with a corresponding nested layer, denoted by $\mathbf{K}_{i}(\cdot)$. The output of $\mathbf{K}_{i}(\cdot)$, represented as $y_{i}$, is computed as follows:
@@
where $x_{i}$ is first added to the output of $\mathbf{K}_{i-1}(\cdot)$, and the resulting feature map is then fed into $\mathbf{K}_{i}(\cdot)$ for further processing. All $y_{i}$ features are concatenated along the channel dimension. Due to the combinatorial explosion effect [18], the output features encapsulate a fusion of receptive field characteristics across different scales and frame levels. These features are then pooled along the time axis to convert frame-level features into utterance-level representations, which are subsequently used to compute the final classification score.
@@ III-A2 Nested Layer
The nested layer acts as the core module responsible for processing the outer layer’s intermediate features, denoted by $x^{\prime}_{i}$, where $i\in\{2,\ldots,s_{1}\}$. Based on Eq. 1, $x^{\prime}_{i}$ is defined as:
@@
Each nested layer $\mathbf{K}_{i}(\cdot)$ is designed to extract multi-scale representations from its input while maintaining computational efficiency. As shown in Fig. 1, the structure of $\mathbf{K}_{i}(\cdot)$ follows a SE-Res2Net-like design, but its input is the feature subset $x^{\prime}_{i}$ from the outer layer of Nes2Net. Specifically, each nested layer consists of the following components:
@@
Multi-Scale Feature Extraction: To enable multi-scale processing, the input feature map $x^{\prime}_{i}$ is equally split into $s_{2}$ subsets along the channel dimension, denoted by $x^{\prime}_{i,j}$, where $j\in\{1,2,\ldots,s_{2}\}$. Each subset undergoes separate transformations through convolutional operations $\mathbf{M}_{j}$ with varying receptive fields, yielding $y_{i,j}$, formulated as:
${y}_{i,j}=\begin{cases}{x^{\prime}}_{i,j}&j=1;\\ \mathbf{M}_{j}({x^{\prime}}_{i,j}+{y}_{i,j-1})&1<j\leq s_{2}.\end{cases}$ (3)
@@
Residual Connections: To enhance gradient flow and stabilize training, a residual connection is applied by adding the input of $x^{\prime}_{i}$ to its output $y_{i}$. This design preserves the original information while incorporating newly learned features.
In summary, the nested layer is lightweight, highly efficient, and designed to improve robustness and generalization across diverse conditions.
@@ III-B Enhanced Nested Res2Net (Nes2Net-X)
Nes2Net efficiently addresses the high-dimensional feature issue. However, it relies on an additive combination method within the nested layer, which may limit the flexibility and effectiveness of feature extraction, as it implicitly assigns equal importance to all features. To further enhance the representational capacity of Nes2Net, we propose an improved variant named Nes2Net-X. It replaces the original addition operation in the nested layer with a concatenation followed by a learnable weighted summation. This design explicitly preserves feature subset individuality before fusion and employs learnable weights to adaptively combine these subsets. The Nes2Net-X consists of the following components:
Feature Splitting and Processing: This component is the same as that in Nes2Net nested layer. The input feature $x^{\prime}_{i}$ is equally split into $s_{2}$ subsets along the channel dimension, denoted by $x^{\prime}_{i,j}$, where $j\in\{1,2,\ldots,s_{2}\}$. Each subset $x^{\prime}_{i,j}$ undergoes a convolutional operation to extract feature representations.
Feature Concatenation: The outputs of the convolutional layers are denoted as $z_{i,j}$. In Nes2Net-X, instead of summing the processed features as in the Nes2Net, each current subset $x^{\prime}_{i,j}$ is concatenated with the previous output $z_{i,j-1}$ along a newly introduced dimension before being processed.
$y_{i,j}=\sum_{k=1}^{s}w_{i,j,k}\cdot z_{i,j,k}$ (4)
where $s$ denotes the number of subsets, $w_{i,j,k}$ represents the weight for the $k$-th subset features $z_{i,j,k}$.
The weighted summation provides more flexible and effective feature integration, offering several advantages: